### v1

# Hypercube Policy Regularization Framework for Offline Reinforcement Learning

Abstract
Offline reinforcement learning has received extensive attention from scholars because it avoids the interaction between the agent and the environment by learning a policy through a static dataset. However, general reinforcement learning methods cannot get satisfactory results in offline reinforcement learning due to the out-of-distribution state actions that the dataset cannot cover during training. To solve this problem, the policy regularization method that tries to directly clone policies used in static datasets has received numerous studies due to its simplicity and effectiveness. However, policy constraint methods make the agent choose the corresponding actions in the static dataset. This type of constraint is usually over-conservative, which results in suboptimal policies, especially in low-quality static datasets. In this paper, a hypercube policy regularization framework is proposed, this method alleviates the constraints of policy constraint methods by allowing the agent to explore the actions corresponding to similar states in the static dataset, which increases the effectiveness of algorithms in low-quality datasets. It was also theoretically demonstrated that the hypercube policy regularization framework can effectively improve the performance of original algorithms. In addition, the hypercube policy regularization framework is combined with TD3-BC and Diffusion-QL for experiments on D4RL datasets which are called TD3-BC-C and Diffusion-QL-C. The experimental results of the score demonstrate that TD3-BC-C and Diffusion-QL-C perform better than state-of-the-art algorithms like IQL, CQL, TD3-BC and Diffusion-QL in most D4RL environments in approximate time.
keywords:
reinforcement learning , Offline reinforcement learning , policy regularization
journal: Neural Networks
[…]
[university]organization=Sun Yat-Sen University,addressline=135 Xingang West Road, Guangzhou, city=Guangzhou, postcode=510275, state=Guandong, country=China
1 Introduction
The objective of Reinforcement Learning(RL) is to learn an optimal policy through the interaction between the agent and the environment. Given the similarity between the manner of RL to generate diverse policies and cognitive abilities, it has received extensive attention and research. At present, there have been some notable achievements in the field, such as games[23, 27], medicine[32],sport[5] and navigator[12].
However, a potential issue with general RL is that the agent must interact with the environment during training, which may cause harm to the environment or the agent, leading to economic losses[3]. One potential solution to this problem is offline Reinforcement learning (offline RL)[19]. In contrast to RL, in offline RL, the agent no longer learns a policy through interaction with the environment. Instead, the agent tries to learn an optimal policy from a static dataset. Since the training of offline RL only requires a static dataset, it does not have the problem of the environment or agent being damaged during training. Nevertheless, offline RL only learns from a limited static dataset, which is typically unable to cover the full spectrum of potential state-action combinations. These state-action pairs that are not included in the static dataset are called out-of-distribution (OOD) state actions[8]. Function estimators employed during offline RL training are prone to inaccuracy when estimating the state-action value (Q-value) or state value (V-value) of an OOD state action. This erroneous overestimation of some Q-values or V-values is likely to result in the agent selecting some suboptimal state-actions during training, This preference for suboptimal state-actions ultimately leads to the inability to obtain a satisfactory policy. To address this issue, offline RL has developed a set of techniques that are distinct from those employed in general RL.
Current offline RL methods can be mainly classified into policy regularization[21, 31], Q-value regularization[1, 22]. importance sampling[11, 34], uncertainty estimation[30, 4] and model-based methods[13, 14, 20, 33]. Among the aforementioned methods, Q-value regularization methods constrain the Q-value of OOD state-actions, thereby obtaining a policy through the exploration of the constrained state action. Policy regularization methods attempt to clone the policy used for static datasets. Current algorithms utilizing policy regularization methods can achieve good results by using it in conjunction with advanced policy clone models such as diffusion models to clone the behavior of policy distribution[28]. Current Q-value regularization methods use ensemble[4], count-based[14], and other methods to more accurately constrain the Q-value of OOD state actions, thereby facilitating enhanced exploration.
The motivation behind this article is as follows:
(1) Policy regularization: One advantage of policy regularization methods is that they have the potential to get a good policy in a shorter time than other methods[7]. However, when faced with low-quality dataset, policy regularization methods will also endeavor to replicate the bad actions within it, resulting in a suboptimal policy[18]. A potential solution to this problem may be found by undertaking an examination of some relevant actions.
(2)Q-value regularization: Q-value regularization methods can perform well in low-quality datasets by exploring the whole space. However, if the agent is allowed to explore the whole space, it is necessary to constrain the Q-value of OOD state-actions, which will result in a significant increase in the number of hyperparameters and operation time. A potential solution to this problem may be found in the reduction of the exploration range of the agent.
(3)Improvement idea: Due to the aforementioned issues, the combination of policy regularization and Q-value regularization methods may yield superior results. Therefore, this article seeks to permit the agent to explore some promising actions and replicate the overall state actions that have been explored. It can be noted that for any given state, there may be multiple similar states within the static dataset. Furthermore, in datasets, especially low-quality datasets, such similar states may have better actions due to the randomness of the sampling actions. Therefore, an attempt is made to set the exploration scope of the agent in a given state limited to actions that correspond to adjacent states.
In light of the above ideas, motivated by Latin hypercube sampling[24], which transforms the space into hypercubes, the hypercube policy regularization framework is proposed. In this framework, the state space is divided into hypercubes and the agent is permitted to explore the actions corresponding to the states in the static dataset within the same hypercube. Compared with current policy regularization methods, which are ineffective in low-quality datasets, and the current Q-constraint methods, which require a long training time[17]. The hypercube policy regularization framework enables agents to explore actions to a certain extent while maintaining sufficient constraints on out-of-distribution state actions, thereby improving the performance of algorithms while maintaining the same training time as policy regularization methods. Another advantage of this framework is its versatility, which allows it to be applied to a variety of policy regularization algorithms such as TD3-BC[7] and Diffusion-QL[28]. The combined algorithm, TD3-BC-C and Diffusion-QL-C exhibited remarkable performance in experiments conducted on the D4RL dataset[6]. TD3-BC-C and Diffusion-QL-C can achieve better results in a similar time compared to those of state-of-the-art(SOTA) algorithms such as Diffusion-QL and IQL[15].
The remainder of this paper is organized as follows. Firstly, the related work and the required basic knowledge of this article are introduced. Secondly, a detailed explanation of the hypercube policy regularization framework is provided. Finally, experiments for the new algorithms are conducted on the D4RL dataset.
2 Related Work
This article is mainly related to policy regularization. In offline RL, The basic idea of the policy regularization method is to solve the OOD state action problem by penalizing the state action deviation from the static dataset with the distance between the trained action state and the corresponding state action within the static dataset. Among the policy constraint methods, BCQ[8] considers the extrapolation error and adds disturbances based on the use of VAE to generate actions to solve the extrapolation error. Therefore, BCQ can directly learn a good policy from a static dataset. BEAR[16] considers that previous algorithms cannot fit out-of-distribution(OOD) state-actions well. To deal with it, BEAR adds weighted behavior clone loss in the optimization process of the policy to solve the problem of OOD state actions by minimizing Maximum Mean Discrepancy(MMD)[10]. Compared with BCQ and BEAR, which require more operations, TD3-BC[7] achieves better results by adding only a few modifications to BC. TD3+BC adds behavior clone regularization terms to the value function and normalizes the state, using techniques similar to BEAR to achieve better results and less training time. BRAC[29] evaluates various methods of behavior clone regularization and proposes to use value penalty and policy regularization to improve algorithm efficiency. Compared with the above algorithms, IQL [15] additionally introduces the state value function to update the Q-value function and policy. Diffusion-QL[28] achieves a state-of-the-art(SOTA) level by using conditional diffusion models for behavior clones and policy regularization. These methods have achieved good results by modifying the behavior cloning model or adding regularization terms. However, as mentioned before, policy regularization methods rely on behavior clones of data within static datasets, which may result in suboptimal results when the quality of static datasets is poor.
On the other hand, advanced algorithms constrain the Q-values of OOD state-actions and thus allow the agent to explore the entire state-action space. For instance, CQL[17], EDAC[2] and PBRL [4] can also have good results in low-quality datasets, but such algorithms usually require large computing time and computing space to find an excellent policy, and such algorithms cannot solve hard tasks well. Considering the characteristics of the distinguishing features of the two methods, local space exploration is integrated into the overall policy constraints. The goal is to enhance the performance of policy regularization methods in low-quality datasets while maintaining the capabilities of policy regularization methods.
3 Preliminaries
Offline RL: In general reinforcement learning, the environment is defined by a Markov decision process as $(S,A,r,P,{\rho_{0}},\gamma)$[25]. Here, $S$ is the state space, $A$ is the action space, $r:S\times A\to[R_{\min},{R_{\max}}]$ is the reward function, $P:S\times A\to S^{\prime}$ is the transition dynamics. By $P$, the agent translates to a new state after the current state, $\rho_{0}\in S$ is the initial state distribution and $\gamma\in(0,1]$ is the discount factor, which is used to adjust the impact of state-actions over time. The goal of RL is to learn a policy $\pi:S\to A$, by maximizing the expected return:
${E_{a_{i}\sim\pi}[\sum_{i=0}^{T-1}\gamma^{i}r(s_{i},a_{i})]}$
over a finite time horizon $T$. In offline RL, the environment is replaced by a static dataset defined as $D=\{(s,a,r,s^{\prime})\}$, here $s$ are the states, $a$ are the actions and $s^{\prime}$ are the next state after actions collected by a policy $\pi$ and the agent no longer interact with the environment to learn a policy but completely learns from the static dataset. The learning objectives and other parts are similar to general RL.
Policy constraint method: In offline RL, the policy constraint method represents a significant approach, This method obtains a policy by constraining the learned policy to align with specific characteristics of the behavior policy. The most fundamental policy constraint method directly obtains a policy through the least squares estimation, it can be expressed as follows:
$\pi=arg{\min}E_{({s_{i}},{a_{i}})\sim D}[\sum_{i=0}^{T-1}(\pi_{\theta}(s_{i})-a_{i})^{2}]$ (1)
Where $\pi_{\theta}$ is the training policy determined by parameter $\theta$, and $a_{i}$ is the action that follows this policy distribution[26]. The above equation directly implements behavior cloning by minimizing the difference between the policy-selected actions and the actions in the static dataset. Subsequent policy regularization methods use different methods to better clone the policy.
4 Proposed Method
In this section, the hypercube policy regularization framework is proposed and a theoretical analysis of the conditions under which the framework can ensure an enhancement in the performance of algorithms is offered. Finally, a new algorithm is created by combining the hypercube policy regularization framework with a baseline offline RL algorithm.
4.1 Hypercube policy regularization framework
The Q-function approximator is capable of accurately estimating the Q-value of state actions $(s,a)$ in static datasets. It is therefore proposed that the Q-function approximator can also be used to accurately estimate the Q-value of $(s^{\prime},a)$ with $|s-s^{\prime}|\leq\varepsilon$. Given these considerations, the hypercube policy regularization framework is proposed.
Figure 1: Hypercubed state space of dimension 3, choosing $\delta=2$.
The hypercube policy regularization framework is a method within the domain of offline RL that enables the agent to explore a range of actions within a hypercube, subsequently utilizing the best actions within this hypercube for policy clone. Compared with general policy regularization methods that only use the corresponding actions in the static dataset for policy clone during training, the hypercube policy regularization framework allows the agent to explore actions in the dataset that correspond to some similar states during the training process. Subsequently, the best action can be employed for policy clones, thereby enhancing the performance of the algorithm.
The hypercube partitioning can be obtained through the following two steps. The initial step is to divide the state space in the static dataset into a hypercube using an integer $\delta$. This is achieved by obtaining the coordinates of the $i$-th dimension of any state $s$ in the static dataset in hypercube space, as outlined below:
$s^{\prime}_{i}=({\delta}({s_{i}}-s_{i}^{\min}))\bmod(s_{i}^{\max}-s_{i}^{\min}),i=1,2,...,n_{1}-1$ (2)
Where $\delta$ determines the accuracy of the hypercube segmentation, $s_{i}^{\min}$ and $s_{i}^{\max}$ are respectively the minimum and maximum values of the $i$-th dimension of states in the static dataset. $n_{1}$ is the dimension of state space.
Second, for the sake of convenience, the coordinates of each state are converted into an integer value:
$v(s^{\prime},a^{\prime})=\sum^{n_{1}-1}_{i=0}{s^{\prime}_{i}\times(\delta+1)}$ (3)
If any two states $s_{i}$ and $s_{j}$ correspond to the same $v(s^{\prime},a^{\prime})$, then the two states are in the same hypercube.
Figure 1 shows a hypercubed state space of dimension 3, the $\delta=2$. Each circle represents a distinct state within the static dataset. The number of states within each hypercube $\geq 0$. During the training phase, the agent is permitted to explore the actions corresponding to the states in the same cube.
After obtaining the above hypercube, for all actions $a_{1},a_{2},...,a_{m}$ in a certain hypercube $v(s^{\prime},a^{\prime})$, the best action $a_{\max}$ can be found by doing the following methodology:
${a_{\max}}=\arg_{a={a_{1}},{a_{2}},...,{a_{m}}}\max Q(s,a)$ (4)
Figure 2: Application of hypercube constraint framework in agent training.
In detail, in any state, the state-action value function $Q(s,a_{\max})$ of selecting the current optimal action within the same hypercube is compared with the state-action value function $Q(s,a)$ of selecting the action corresponding to that state itself, in order to select the better action and achieve the goal of enabling the agent to explore within a certain range. Finally, the hypercube policy regularization framework can be derived by integrating the constraint with a Markov decision process. Figure 2 illustrates a training flow chart that excludes the evaluation phase. At the beginning of training, the static dataset is hypercubed and it was used to constrain the current action $a$ to obtain a new action $a_{new}$ during training. At the end of a loop, the process is reversed to update the policy $\pi_{\theta}$, and this step is repeated to output a policy $\pi_{\theta}$.
4.2 Superiority analysis
If the Q-function approximator provides an entirely correct estimate of the Q-value, then it follows that utilizing the hypercube policy regularization framework can maintain or enhance the performance of the algorithm under any circumstances. However, in practical applications, the issue of imprecise estimation of the Q-value must be taken into account. When one state in a hypercube takes the optimal action corresponding to another state within the same hypercube, the Q-value corresponding to the state action $(s,a)$ may be overestimated. It is demonstrated that modifying the value of $\delta$ allows the hypercube constraint framework to enhance or sustain the performance of the algorithm regardless of the precision of the Q-value estimation.
Assumption 1.
Assuming that the Q-function learned satisfies the Lipschitz condition, that is there exists a positive constant $K$ for all $(s_{1},a_{1}),(s_{2},a_{2})$:
$\left|{Q({s_{1}},{a_{1}})-Q({s_{2}},{a_{2}})}\right|\leq K\left|{({s_{1}}\oplus{a_{1}})-({s_{2}}\oplus{a_{2}})}\right|$ (5)
Where $\oplus$ is the vector concatenation operation, this assumption can be satisfied when the Q-function is approximated by neural networks[9]. Thus, we demonstrate that the hypercube policy regularization framework ensures that the algorithm’s performance does not deteriorate under Assumption 1.
Assumption 2.
Assuming that the Q-function approximator can accurately estimate the Q-value of state-actions within the static dataset.
Since the Q-function approximator estimates the Q-values of all state-actions through the data in the static dataset, It seems reasonable to posit that the Q-function approximation is capable of providing an accurate estimate of the Q-value of state-actions within the static dataset.
Theorem 1.
When choosing an appropriate hypercube segment precision $\delta$ within the hypercube policy regularization framework, the performance of the algorithms is not diminished for any state $s$, that is:
$Q(s,{a^{\pi_{new}}})\geq Q(s,{a^{\pi_{old}}})$ (6)
Proof.
When only the $(s,a)$ exists in the hypercube, then $\pi_{new}=\pi_{old}$ and $Q(s,a^{\pi_{old}})=Q(s,a^{\pi_{new}})$.
When there is another state-action $(s^{\prime},a^{\prime})$ in the hypercube, according to Assumption 2, there is an accurate estimate of the Q-value relationship between $(s,a)$ and $(s^{\prime},a^{\prime})$. Therefore, $Q(s,a^{\prime})$ is discussed through $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$. For this, three cases must be considered: $Q(s,a)<Q(s^{\prime},a^{\prime})$, $Q(s,a)=Q(s^{\prime},a^{\prime})$ and $Q(s,a)>Q(s^{\prime},a^{\prime})$.
When $Q(s,a)>Q(s^{\prime},a^{\prime})$, $Q(s,a)$ can be expressed as:
$Q(s,a)=Q(s^{\prime},a^{\prime})+b_{1}$ (7)
Here, $b_{1}>0$. By Assumption 1, the following inequality can be obtained:
$\begin{array}[]{c}Q(s,a^{\prime})\leq Q(s^{\prime},a^{\prime})+K|(s\oplus a^{\prime})-(s^{\prime}\oplus a)|\\ =Q(s^{\prime},a^{\prime})+K|s-s^{\prime}|\end{array}$ (8)
combing Equation(7) and inEquation(8), it can be known that when $K|s-s^{\prime}|<b_{1}$ , $Q(s,a^{\prime})\leq Q(s,a)$.
In a hypercube, the longest distance between two state-actions is $S_{max}$. $S_{max}\geq|s-s^{\prime}|$ and $S_{max}$ is bounded due to the limited data in the static dataset. Because the longest distance in the hypercube is proportional to the size of the hypercube. When $\theta$ increases, $S_{max}$ decreases proportionally. Therefore, there exists an integer $c_{1}$, when $\theta>c_{1}$,
$K|s-s^{\prime}|\leq KS_{max}\leq b_{1}$ (9)
This guarantees that when $\theta>c_{1}$, $Q(s,a)>Q(s^{\prime},a^{\prime})$, the new policy is always worse than the original policy, so $\pi_{new}=\pi_{old}$, $Q(s,a^{\pi_{new}})=Q(s,a^{\pi_{old}})$ in this situation.
When $Q(s,a)=Q(s^{\prime},a^{\prime})$, $\pi_{new}=\pi_{old}$ is selected to obtain a conservative policy. Therefore, in this case, $Q(s,a^{\pi_{new}})=Q(s,a^{\pi_{old}})$.
While $Q(s^{\prime},a^{\prime})>Q(s,a)$, $Q(s^{\prime},a^{\prime})$ can be expressed as:
$Q(s^{\prime},a^{\prime})=Q(s,a)+b_{2}$ (10)
Here, $b_{2}>0$. By Assumption 1, the following conclusions can be drawn:
$\begin{array}[]{c}Q(s,a^{\prime})\geq Q(s^{\prime},a^{\prime})-K|(s\oplus a^{\prime})-(s^{\prime}\oplus a^{\prime})|\\ =Q(s^{\prime},a^{\prime})-K|s-s^{\prime}|\end{array}$ (11)
Combining Equation(10) and inEquation(11), it can be known that when $K|s-s^{\prime}|<b_{2}$, $Q(s,a^{\prime})\geq Q(s,a)$. Similar to $Q(s^{\prime},a^{\prime})<Q(s,a)$, there exists an integer $c_{2}$, when $\theta>c_{2}$,
$K|s-s^{\prime}|\leq KS_{max}\leq b_{2}$ (12)
This demonstrates that when $\theta>c_{2}$, $Q(s^{\prime},a^{\prime})>Q(s,a)$. Therefore, regarded $a^{\prime}$ as $a_{new}$, the new policy can guarantee that $Q(s,a^{\pi_{new}})\geq Q(s,a^{\pi_{old}})$.
For the situation that there is only one extra $(s^{\prime},a^{\prime})$ in the hypercube and $\theta>max\{c_{1},c_{2}\}$. When $Q(s^{\prime},a^{\prime})\leq Q(s,a)$, the policy remains unchanged, $Q(s,a^{\pi_{new}})=Q(s,a^{\pi_{old}})$. When $Q(s^{\prime},a^{\prime})>Q(s,a^{\prime})$, $Q(s,a^{\pi_{new}})\geq Q(s,a^{\pi_{old}})$. Therefore, inEquation(6) holds in this case.
When there are multiple state-actions $(s_{i},a_{i}),i=1,2,...,n$ within the static dataset in each hypercube, refer to the above analysis of the existence of only one additional state-action in the hypercube, it can be seen that When the Q-value of some state-actions is smaller than $Q(s,a)$, there exist $b_{1}>0$, $\theta>c_{1}$ make:
$Q(s,a)\geq Q(s,a_{i}),i=1,2,..,n$ (13)
While $K|s-s_{i}|\leq KS_{max}\leq b_{1}$, $Q(s,a)\geq Q(s_{i},a_{i})$
When the Q-value of some state-actions is larger than $Q(s,a)$, there also exist $b_{2}>0$ and $\theta>c_{2}$ make:
$Q(s,a)\leq Q(s,a_{i}),i=1,2,..,n$ (14)
While $K|s-s_{i}|\leq KS_{max}\leq b_{2}$, $Q(s,a)\leq Q(s_{i},a_{i})$.
Therefore, when there are multiple state-actions in a hypercube, selecting $\theta>max\{c_{1},c_{2}\}$ can also make $Q(s,{a^{\pi_{new}}})\geq Q(s,{a^{\pi_{old}}})$.
This means that the hypercube policy regularization framework guarantees that the performance of the algorithm is not diminished, thereby enabling the framework to be applied in many scenarios.
In theory, selecting a large value for the parameter $\delta$ can ensure that the performance of the algorithms remains unaltered or improves. In practice, the Q-function approximator is capable of producing a relatively accurate estimate of the explored state actions within the static dataset, Consequently, the uncertainty associated with the Q-value obtained from exploring actions within the static dataset is likely to be relatively low. Therefore, to improve the effectiveness of the algorithms, smaller $\delta$ should be used to make the agent explore more actions so that the agent can maintain a balance between conservatism and aggressiveness and thus achieve the best results.
4.3 Algorithms
Algorithm 1 TD3-BC-C
Input: Initial policy parameter ${\it\theta}$; Target policy parameter $\theta^{-}$; min-batch size $k$; Q-function parameters $\phi_{1}$,$\phi_{2}$; Target Q-function parameters $\phi_{1}^{-}$, $\phi_{2}^{-}$; offline dataset ${\it D}$; actions obtained by hypercube policy regularization ${a^{\prime}}$
Output: Policy $\pi_{\theta}$
[…]
3:        Sample a mini-batch of $k$ state-action pairs $\left\{(s_{t},a_{t}),t=1,..,k\right\}$ from ${D}$
4:        Sample $a_{t+1}\sim\pi_{\theta}(a_{t+1}|s_{t+1})$
5:        Update $Q_{{\phi_{1}}}$ and $Q_{{\phi_{2}}}$ by $\begin{array}[]{l}E_{({s_{t}},a_{t},s_{t+1})\sim D,a_{t+1}\sim{\pi_{\theta}}}[(r({s_{t}},{a_{t}})\\ +\gamma\min Q_{{\phi_{i}}}({s_{t+1}},{a_{t+1}})-Q_{\phi_{i}}({s_{t}},{a_{t}}))^{2}]\end{array}$
6:        ##\## Policy learning
7:        Compared $\mathop{\min}\limits_{i=1,2}{Q_{{\phi_{i}}}}(s,a)$ and $\mathop{\min}\limits_{i=1,2}{Q_{{\phi_{i}}}}(s,a^{\prime})$ to Update ${a^{\prime}}$
8:        Update policy by Equation(15)
9:        # Update target network
10:        Update policy target by ${\theta^{-}}=\rho{\theta^{-}}+(1-\rho)\theta$
11:        Update Q target by: $\phi_{i}^{-}=\rho\phi_{i}^{-}+(1-\rho)\phi_{i}$
12:     end while
13:  end while
This section outlines how the hypercube policy regularization framework can be integrated with offline RL. The hypercube constraint framework can be directly applied to policy regularization algorithms such as TD3-BC and Diffusion-QL.
Taking TD3-BC as an example, our algorithm made only two straightforward modifications to TD3-BC. Firstly, a hypercube was constructed using Equation(3) and employed to store the optimal action within the hypercube. Each time the state $s$ corresponding to the hypercube is selected, a comparison is made between $Q(s,a)$ and $Q(s,a^{\prime})$ to determine whether action $a$ should be replaced within the hypercube. Secondly, in policy updates, the action $a_{h}$ from the hypercube is employed in place of $a$ from the static dataset for policy regularization. The policy update equation is given by:
$\arg\min L(\theta)={L_{c}}(\theta)+{L_{q}}(\theta)$ (15)
Where, $\theta$ is the policy parameter, and $L_{c}$ is the policy regularization loss function.
In TD3-BC and other policy regularization algorithms that try to clone the policy in the static dataset, the $L_{c}$ is:
$E_{(s,a)\sim D}[{(a-\pi_{\theta}(s))^{2}}]$ (16)
Unlike the $L_{c}$ above, the $L_{c}$ is:
$E_{(s,a)\sim D}[{(a_{h}-\pi_{\theta}(s))^{2}}]$ (17)
Where $a_{h}$ represents the action within the static dataset that has the highest Q within the hypercube. Given that $Q(s,a_{h})\geq Q(s,a)$ and Theorem 1, it follows that the actions $a_{h}$ within the hypercube will produce results that are either equal to or better than those produced by $a$ within the static dataset. Therefore, the utilization of Equation(3) for behavior cloning can lead to a more effective policy.
The regularization term, $L_{q}$ varies across different algorithms. In TD3-BC, the $L_{q}$ is as follows:
$L_{q}=\phi E_{(s_{i},a_{i})\sim D}[\sum_{i=0}^{T-1}Q(s_{i},a_{i})]/E_{(s_{i},a_{i})\sim D}[\sum_{i=0}^{T-1}\left|Q(s_{i},a_{i})\right|]$ (18)
Here $\phi$ is the hyperparameter.
Algorithm 1 can be obtained by regularizing the policy as above.
In pseudocode, similar to TD3-BC, Diffusion-QL differs from TD3-BC only in $L_{q}$, which is:
$L_{q}=\gamma{E_{s_{i}\sim D,a_{i}^{\prime}\sim{\pi_{\theta}}}}[\sum_{i=0}^{T-1}{Q_{\phi}}(s_{i},a_{i}^{\prime})]$ (19)
Here $\gamma=\eta/{E_{(s_{i},a_{i})\sim D}}[\sum_{i=0}^{T-1}|{Q_{\phi}}(s,a)|]$, $\eta$ is a hyperparameter.
Table 1: In the table, half denotes halfcheetah; hp denotes hopper; w denotes walker2d; r denotes random; m denotes medium; and e denotes expert, with the optimal result in each environment is highlighted in bold.
Env IQL CQL TD3-BC Diffusion-QL TD3-BC-C Diffusion-C
half-r $11.0{\pm}2.5$ $\textbf{28.3}{\pm}\textbf{0.5}$ $11.7{\pm}0.7$ $21.0{\pm}0.6$ $26.5{\pm}1.6$ $23.8{\pm}0.6$
hp-r $7.8{\pm}0.4$ $16.4{\pm}14.5$ $8.5{\pm}0.2$ $8.1{\pm}0.1$ $\textbf{31.3}{\pm}\textbf{0.0}$ $13.1{\pm}7.8$
w-r $6.4{\pm}0.1$ $4.2{\pm}1.4$ $1.5{\pm}0.7$ $3.4{\pm}2.0$ $0.0{\pm}0$ 0.0 $\textbf{20.2}{\pm}\textbf{2.9}$
half-m $47.7{\pm}0.2$ $54.9{\pm}0.2$ $48.1{\pm}0.1$ $50.0{\pm}0.2$ $\textbf{64.8}{\pm}\textbf{0.8}$ $54.0{\pm}0.0$
hp-m $55.4{\pm}2.9$ $42.7{\pm}37.6$ $57.4{\pm}2.3$ $74.3{\pm}6.0$ $\textbf{100.6}{\pm}\textbf{1.0}$ $97.9{\pm}4.0$
w-m $77.5\pm 2.3$ $73.6\pm 2.0$ $83.7\pm 1.2$ $86.3\pm 0.4$ $\textbf{93.8}\pm\textbf{1.0}$ $84.6\pm 0.3$
half-m-r $43.7\pm 0.3$ $51.8\pm 0.0$ $45.1\pm 9.8$ $46.9\pm 0.6$ $\textbf{53.6}\pm\textbf{1.0}$ $50.1\pm 0.4$
hp-m-r $55.4\pm 0.4$ $94.4\pm 4.0$ $62.7\pm 20.9$ $\textbf{100.7}\pm\textbf{0.5}$ $99.8\pm 0.8$ $98.8\pm 0.6$
w-m-r $69.3\pm 5.8$ $84.0\pm 3.6$ $76.3\pm 4.1$ $\textbf{93.8}\pm\textbf{0.5}$ $86.3\pm 1.6$ $\textbf{93.7}\pm\textbf{2.7}$
half-m-e $86.7\pm 4.4$ $80.4\pm 12.7$ $93.2\pm 1.9$ $\textbf{96.0}\pm\textbf{0.1}$ $92.5\pm 0.7$ $\textbf{96.1}\pm\textbf{0.6}$
hp-m-e $44.2\pm 4.3$ $99.2\pm 6.1$ $96.8\pm 3.8$ $107.6\pm 4.0$ $\textbf{109.3}\pm\textbf{5.2}$ $105.8\pm 0.9$
w-m-e $109.3\pm 1.0$ $101.2\pm 17.6$ $110.2\pm 0.0$ $109.9\pm 0.2$ $\textbf{115.2}\pm\textbf{1.7}$ $109.8\pm 0.0$
Average $51.2\pm 2.1$ $60.9\pm 8.4$ $58.0\pm 4.4$ $66.4\pm 1.2$ $\textbf{72.7}\pm\textbf{1.2}$ $70.3\pm 1.8$
5 Experiment
The objective of the experiments presented in this section is to validate the hypothesis that the utilization of the hypercube policy regularization framework can enhance the performance of policy regularisation algorithms in offline RL. To this end, our algorithm was validated using the D4RL dataset, a popular benchmark for this field, and the results were compared with those obtained by other well-known algorithms.
A variety of tasks were considered, drawn from a range of domains within the D4RL benchmark, including Gym, AntMaze and Adroit.
5.1 Experiment in Gym
5.1.1 Experiment setting
Gym datasets are currently the most commonly used evaluation criterion tasks. Gym datasets typically have high-quality training curves, and the reward functions in Gym datasets are relatively smooth. Therefore, Gym datasets are relatively simple compared with other tasks.
To verify the efficacy of the hypercube policy regularization framework, it was integrated with TD3-BC and Diffusion-QL to create the TD3-BC-C and Diffusion-QL-C algorithms. The algorithms were then tested on the D4RL dataset, and the test results were compared with some classical and advanced algorithms. The compared algorithms include TD3-BC[7], which is one of the benchmark algorithms; the Diffusion-QL algorithm[28], which is the sota policy constraint algorithm and constitutes another benchmark algorithm; IQL[15], which is a policy constraint algorithm that uses a combination of V-value and Q-value functions; CQL[17], a classical Q-value constraint algorithm.
Training is performed for 2000 epochs in the medium-expert environment and 1000 epochs for other types of environments, 5 random seeds are used for each environment. The results of the training can be seen in Table LABEL:result, and the training curves can be seen in FigureLABEL:exp.
5.1.2 Experiment result
Table 2: In the table, max represents the best result obtained during training and the corresponding error at this time, the optimal result in each environment is highlighted in bold.
        Env         TD3-BC         Diffusion-QL         TD3-BC-C         Diffusion-C
        maze-u         $47.2{\pm}47.1$         $45.8{\pm}59.0$         $\textbf{95.9}{\pm}\textbf{8.9}$         $91.0{\pm}4.8$
        maze-u-d         $40.3{\pm}8.9$         $56.8{\pm}32.5$         $59.3\pm 3.0$         $\textbf{70.9}{\pm}\textbf{18.7}$
        maze-m-d         $3.2{\pm}2.5$         $0.0{\pm}0$ 0.0         $\textbf{62.1}{\pm}\textbf{5.1}$         $0.0{\pm}0$ 0.0
        maze-m-p         $0.4{\pm}0.6$         $0.0{\pm}0$ 0.0         $\textbf{59.9}{\pm}\textbf{8.7}$         $0.4{\pm}3.5$
        maze-l-d         $0.0{\pm}0$ 0.0         $14.2{\pm}29.2$         $25.2{\pm}6.7$         $\textbf{31.6}{\pm}\textbf{32.6}$
        maze-l-p         $0.0\pm 0$ 0.0         $2.2\pm 5.6$         $20.1\pm 6.7$         $\textbf{25.5}\pm\textbf{14.9}$
        maze-u-max         $95.3{\pm}1.2$         $90.2{\pm}2.7$         $94.0{\pm}1.9$         $\textbf{95.9}{\pm}\textbf{8.9}$
        maze-u-d-max         $70.5{\pm}7.6$         $\textbf{75.7}{\pm}\textbf{7.1}$         $59.3{\pm}3.0$         $75.4{\pm}9.9$
        maze-m-d-max         $6.0{\pm}1.6$         $75.6{\pm}5.3$         $66.4{\pm}4.3$         $\textbf{82.9}{\pm}\textbf{4.8}$
        maze-m-p-max         $2.8{\pm}4.9$         $54.8{\pm}14.7$         $\textbf{59.9}{\pm}\textbf{8.7}$         $46.0{\pm}37.3$
        maze-l-d-max         $0.7{\pm}1.1$         $50.3{\pm}5.7$         $27.7{\pm}10.2$         $\textbf{55.0}{\pm}\textbf{13.7}$
        maze-l-p-max         $0.5\pm 0.8$         $\textbf{50.4}\pm\textbf{4.8}$         $24.8\pm 1.9$         $39.9\pm 11.6$
        pen-c         $5.0\pm 3.7$         $33.0\pm 8.7$         $\textbf{46.0}\pm\textbf{1.3}$         $31.0\pm 8.7$
        pen-h         $-2.4\pm 2.2$         $57.8\pm 8.6$         $\textbf{68.6}\pm\textbf{8.8}$         $49.9\pm 1.1$
        Average         $19.3\pm 5.9$         $43.3\pm 13.1$         $\textbf{54.9}\pm\textbf{5.7}$         $49.7\pm 12.2$
Table LABEL:result demonstrates that TD3-BC-C exhibits optimal performance in $7$ out of $12$ environments. When compared with TD3-BC, it shows better performance in all environments except walked-random. Moreover, the performance of TD3-BC-C is superior to that of the current state-of-the-art algorithm, Diffusion-QL. Diffusion-QL-C displays a markedly enhanced performance in environments of inferior quality such as random environments. Furthermore, the overall performance of Diffusion-QL-C is also improved compared with Diffusion-QL.
It can be seen that although the performance of Diffusion-QL is much better than TD3-BC, the performance of TD3-BC-C is better than Diffusion-QL-C. It is hypothesized that this discrepancy can be attributed to the fact that Diffusion-QL has more hyperparameters compared with TD3-BC. The use of excessive hyperparameters may result in a reduction in the sensitivity of the hypercube policy regularization framework, thereby limiting the efficacy of the hypercube policy regularization framework in Diffusion-QL. Nevertheless, it is hypothesized that Diffusion-QL-C may demonstrate remarkable performance if the hyperparameters are subjected to further adjustment.
5.2 Experiment in other D4RL environment
To further validate the performance of the algorithms, additional experiments were conducted in Adroit and Maze environments. Compared with Gym environments, Adroit and Maze environments are more complex and challenging for agents to learn. Maze environments comprise a series of maze navigation tasks, wherein the agent is required to develop a policy to reach the end point from the path information provided in the static dataset. The agent is rewarded solely upon reaching the designated endpoint. It is therefore necessary for the algorithm to possess the ability to overcome sparse rewards. Adroit environments consist of some datasets that simulate a range of human activities. The complexity of human activities, the limited scope of the data and the relatively sparse rewards make it extremely challenging to learn Adroit environments.
The original reward was utilized for the experiments conducted on Gym and Adroit. For the experiments on Maze, since the majority of the algorithms were unable to achieve effective results with the original reward information. Consequently, a modification was made to this information following algorithms such as Diffusion-QL. In the Maze and Adroit environments, the algorithms employed for comparison are TD3-BC and Diffusion-QL, as the special parameters used by most algorithms in these environments have not been made available for public scrutiny.
In order to ensure a fair comparison, the results presented in Table LABEL:result2 include the results obtained after training for 1000 epochs as well as the highest score achieved during the training process. This is due to the significant performance fluctuations of algorithms in maze environments. The training diagrams are illustrated in FigureLABEL:exp2.
The experimental results demonstrate that TD3-BC-C and Diffusion-C are also effective in complex environments, which further demonstrates the effectiveness of the hypercube policy regularization framework.
5.3 Hyperparameter setting
Table 3: The number of the state space and action space has been partitioned in different D4RL environments and the uncertainty constraint parameter $\delta$.
Environment state_dim $\delta$ epoch
hp-r 3 20 1000
hp-m-r 3 20 1000
hp-m 3 20 1000
hp-m-e 3 20 2000
half-r 6 5 1000
half-m-r 6 5 1000
half-m 6 5 1000
half-m-e 6 5 2000
w-r 6 5 1000
w-m-r 6 5 1000
w-m 6 5 1000
w-m-e 6 5 2000
maze-d 2 5 1000
maze-m 2 5 1000
maze-l 2 5 1000
pen-c 24 5 1000
pen-h 24 5 1000
The hypercube constraint framework introduces only one additional hyperparameter $\delta$, which represents the segmentation accuracy of the state space. Therefore, the value of $\delta$ can be determined based on its practical significance. In particular, the hyperparameter $\delta$ can be selected based on the dimensionality of the state space and the quality of the dataset.
For low-quality datasets, a lower $\delta$ is selected to permit the agent to explore more actions. In contrast, for higher-quality datasets, a higher $\delta$ is chosen, whereas for static datasets, policy cloning is deemed sufficient. As for the dimension of the space, the probability of two state actions occurring in the same hypercube is lower for higher dimensional spaces when $\delta$ remains constant. Consequently, a lower $\delta$ is selected for higher dimensional environments and a higher $\delta$ is selected for lower dimensional environments. In order to ensure a fair comparison, the majority of environments in this article have been configured with identical parameters. The selection of specific parameters in complex Maze and Adroit environments may result in better performance.
Table 3 illustrates the specific $\delta$ that was employed and lists some of the parameters that were deemed to be relevant.
5.4 Computational cost
Table 4: The computational cost of some algorithms in hopper medium. GPU Memory is the average of GPU Memory used in hopper-medium.
Algorithm Runtime(s) GPU Memory(GB)
IQL 14.6 2.0
CQL 62.2 1.4
TD3-BC 13.8 1.9
Diffusion-QL 68.9 2.0
TD3-BC-C 21.0 1.9
Diffusion-QL-C 77.3 2.0
To demonstrate the efficacy of the hypercube policy regularization framework from a different perspective, the runtime and space requirements of each algorithm are compared. The running time and space were obtained by executing hopper-medium for one epoch, the results are presented in Table 4. Table 4 reveals that TD3-BC-C and Diffusion-QL-C only increase a small amount of runtime and GPU Memory compared with TD3-BC and Diffusion-QL. It can therefore be concluded that the hypercube policy regularization framework has significant advantages in terms of performance and computational cost.
6 Conclusions
In this paper, the hypercube policy regularization framework is proposed, a novel framework for policy regularisation. This framework alleviates the constraints of policy regularization methods by enabling the agent to explore actions corresponding to similar states. It is indicated both theoretically and experimentally that the hypercube constraint framework can improve the performance of algorithms with policy regularization. This is achieved by alleviating constraints on OOD state actions and enabling the agent to explore some actions. In D4RL experiments, TD3-BC-C and Diffusion-QL-C obtained by the hypercube policy regularization framework demonstrated favorable results with the introduction of only one additional hyperparameter and a modest increase in computational cost. The performance of the hypercube policy regularization framework demonstrates that further research on the optimal utilization of static datasets has important implications for offline reinforcement learning. Therefore, we intend to pursue further research into the utilization of static datasets.
Appendix A Ablation experiment
As for the sensitivity of the parameter $\delta$, we illustrate that the hyperparameter $\delta$ has a negligible impact on the performance of the algorithm. Experiments are conducted with different $\delta$ in walker2d-medium and hopper-medium-replay environments for testing. The experimental results are shown in Figure LABEL:abl.
TD3-BC-C and Diffusion-QL-C are equivalent to TD3-BC and Diffusion-QL respectively When the value of a is very large. Therefore, TD3-BC and Diffusion-QL are employed to replace the situation when $\delta$ selects a larger value in the ablation experiment.
It can be seen from Figure LABEL:abl, in walker2d-medium-replay and hopper-medium, TD3-BC-C and Diffusion QL-C can both have results when the value of $\delta$ is not too extreme. However, when $\delta=1$, the exploration range of the agent is excessive, and the constraints on the out-of-distribution state actions are insufficient, rendering unsatisfactory results in some environments. When $\delta$ is excessively large, the hypercube policy regularization framework merely clones the behavior of the current state action, yielding only suboptimal outcomes.
Appendix B Parameters
Table 1: Partial hyperparameters in GPC-SAC
Hyperparameter value
discount 0.99
policy_rl 3e-4
qf_rl 3e-4
soft_target_tau 5e-3
train_pre_loop 1000
Q-network FC(256,256,256)
Optimizer Adam
activation_function relu
$\beta$ 1
$\beta_{next}$ 0.1
$\phi$ 2.5,40.0
$\eta$ 1.0
For the hyperparameters utilized in the Gym experiment, since our algorithm has been modified from TD3-BC and Diffusion-QL. Consequently, the majority of our hyperparameters are consistent with those of these two algorithms, as illustrated in Table 1.
In Adroit and Maze environments, refer to the parameter in  [28, 21]. The selection of parameter $\phi$ is $\{0.1,2.5,10,25\}$ for {pen,maze-u,maze-m,maze-l}, parameter $\eta$ as $\{0.1,0.15,0.5,2.0,2.0,3.0,4.5,3.5\}$ for {pen-c,pen-h,maze-u,maze-u-d,maze-m-p,maze-m-d,maze-l-p,maze-l-d}.
Appendix C Implement Detail
Our experimental results were obtained through the following method. The specific implementation of TD3-BC-C, Diffusion-B can be found at https://github.com/lastTarnished/GPC-SAC. Experiments are conducted on D4RL, which can be publicly used at https://github.com/Farama-Foundation/D4RL. The following are the implementations of the comparative algorithms in our experiment. (1)For CQL, https://github.com/lastTarnished/Hypyercud-Policy-Regularization/ has been referenced, which is the official implementation of CQL. (2)For the implementation of IQL, https://github.com/ikostrikov/implicit_q_learning has been referenced, this code has simplified the implementation of IQL compared with the official code, resulting in a slight decrease in performance in a few environments. (3)For the implementation of TD3-BC, https://github.com/sfujim/TD3_BC and no changes were made to the core code. (4)For the implementation of Diffusion-QL, https://github.com/Zhendong-Wang/Diffusion-Policies-for-Offline-RL has been referenced and no changes were made to the core code.